OpenStudy (anonymous):

Maya ran a 100-yd sprint.How many feet did she sprint? 1 yd = 3 ft

OpenStudy (amistre64):

what do you multiply 1 yard by to get 100 yards?

OpenStudy (skullpatrol):

Whenever you see a conversion factor like 1yd = 3ft it can be written as a fraction two ways. For this question you would write it as \[\frac{ 3ft }{ 1yd }\]and multiply 100yds. by the fraction. (note how the unit of yds cancels out leaving the unit of ft in the answer. \[100yds \times \frac{ 3ft }{ 1yd } = 100 \times 3 ft \]

OpenStudy (skullpatrol):